How many kg of dies apricots worth 8.30 per kilogram should be mixed with dried apples worth 6.50 per kilgram to produce 12 kg of a mixture worth 6.95 per kilogram?
Answer by stanbon(75887)

How many kg of dried apricots worth 8.30 per kilogram should be mixed with dried apples worth 6.50 per kilgram to produce 12 kg of a mixture worth 6.95 per kilogram?
---------------
Equations:
Quantity Eq: p + a = 12 kg
Value Eq:: 8.3p+6.5a = 6.95*12
----------------------------------------
Multiply thru Quantity by 830
Multiply thru Value by 100
----------------------------------
830p + 830a = 830*12
830p + 650a = 695*12
--------------------------
180a = 135*12
a = 9 (# of kg of apples)
---
p = 12-9 = 3 kg (# of kg of apricots)
===========================================
Cheers,
Stan H.
==============